OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition: slope 6 and passes through (5,20)

OpenStudy (cwrw238):

use the general form y-y1 = m(x-x1) where m = slope = 6 and (x1,y1) = (5, 20)

OpenStudy (cwrw238):

plug in m=6, x1 = 5 and y1 = 20 into the general equation and simplify

OpenStudy (anonymous):

y-20=6(x-5) y=6(x-5)+20 y = 6x-10